OpenStudy (anonymous):

Evaluate the expression and write the result in the form a + bi. (5 + 3i)(2 − i) over/3 + i

OpenStudy (anonymous):

Solve it and Rationalise the Denominator ... :).

OpenStudy (anonymous):

if i knew how to do that i wouldnt be asking lol

OpenStudy (anonymous):

\[(5+3i)(2-i)=(10+3)+(6-5)i=13+i\]

OpenStudy (anonymous):

so you have \[\frac{13+i}{3+i}=\frac{13+i}{3+i}\times \frac{3-i}{3-i}\]

OpenStudy (anonymous):

the denominator is \[9+1=10\] and that is why you use this method

OpenStudy (anonymous):

the numerator is \[(13+i)(3-i)=40-10i\]

OpenStudy (anonymous):

so you get \[\frac{40-10i}{10}=4-i\]

OpenStudy (anonymous):

so is the final answer 4-i

OpenStudy (anonymous):

o ok good

OpenStudy (anonymous):

that is what i got, yes